OpenStudy (anonymous):

A cone has a base area of 3 in² and a lateral area which is 1.5 times bigger than the base area. The surface area of the cone is _____ in².

OpenStudy (anonymous):

Lateral surface area =3*1.5=4.5in Surface area = lateral area + base area =4.5+3=7.5